Define a function
f
that determines the amount that Esmeralda pays (in dollars) for gas to drive
m
miles, given that Esmeralda's car gets 26 miles per gallon and gas costs $3.70 per gallon

Answer by Theo(13342)

f(m) = m/26 * 3.70

m is the number of miles driven.

for example:

if she drove 260 miles, then the number of gallons she bought was 260 / 26 = 10 and the price she paid was 10 * 3.70 = 37.00 dollars.